OpenStudy (anonymous):

Solve the equation without using logarithms. 3^9x=9^3x-6

OpenStudy (anonymous):

3^9x = (3^2)^3x -6

OpenStudy (anonymous):

3^9x = 3^6x-6 3^9x - 3^6x = -6

OpenStudy (anonymous):

theres one answer x=

OpenStudy (anonymous):

is it 3^(9x)=9^(3x-6)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

-4?

OpenStudy (anonymous):

o...

OpenStudy (anonymous):

3^(9x) = 3^(6x-12) 9x = 6x -12 3x = -12 x = -4

OpenStudy (anonymous):

\[3^{9x}=9^{3x-6}\]

OpenStudy (anonymous):

-4 is correct

OpenStudy (anonymous):

where did you get 6x-12 from?

OpenStudy (anonymous):

\[9^{3x-6}=(3^{2})^{3x-6}=3^{6x-12}\]

OpenStudy (anonymous):

how does (3^2)^3x-6 give you 3^6x-12?

OpenStudy (anonymous):

multiply by 2 obviously but im a little confused

OpenStudy (anonymous):

(a^m)^n=a^(m*n)

OpenStudy (anonymous):

u just multiply the powers together?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok i see... THANKS a lot!